(((-17-1+26-38=-3-m-8+2m)))
Please help me solve this problem
I got:
-17-1+26-38 = -30
-3-m-8+2m=
-11+1m=
This is where I'm totally lost...

Answer by longjonsilver(2297)

You are correct so far:

-17-1+26-38=-3-m-8+2m
-30 = -11+m

Now add 11 to both sides, giving -19 = m

that is your answer... m is -19

jon.